OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. (2 points) Question 2 options: 1) -334 2) -274 3) -284 4) -346

OpenStudy (welshfella):

the common difference = -226 - 26 --------- 21

OpenStudy (welshfella):

nth term = a1 + d(n - 1) where a1 = first term = 26 , n =31 for 31st term and d is calculated as above

OpenStudy (welshfella):

any questions?

OpenStudy (anonymous):

26+-226-26/21*31-1

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

i dont understand this

OpenStudy (welshfella):

nth term = a1 + d(n - 1 31st term = 26 + (-12)(31-1) = 26 - (12*30)

OpenStudy (welshfella):

= 26 - 360 = ?

OpenStudy (anonymous):

So its A?

OpenStudy (anonymous):

-334

OpenStudy (welshfella):

d = -12

OpenStudy (welshfella):

yes A

OpenStudy (anonymous):

thank you :) @welshfella

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (welshfella):

plz create a new question

OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. (2 points) Question 1 options: 1) 203 2) 197 3) 168 4) 163

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

do u goi to conencitons 8grade

OpenStudy (anonymous):

what ? @jjjd

OpenStudy (anonymous):

connections 8grade

OpenStudy (anonymous):

no